# Measuring Compliance of Consent Revocation on the Web

### v2

@@ 5.1.1 Compliant revocation interface implementation (zero or one steps)
Out of the 161 websites, only 9 (5.6%) offer a persistent icon or button floating on the page, thus requiring zero steps to reach the revocation option. $6767$ (41.6%) websites offered a link option in the footer of the page, requiring one step to reach the revocation option. 11 (6.8%) websites showed a consent banner or icon when accessing the privacy policy page from the footer, therefore also requiring one step to revoke consent. All these implementations found on 87 (54%) websites (labeled as Icon, Footer Options and Banner on Policy in Table IV) comply with revocation GDPR principles and consent requirements (LR1-3) since they are presented within the same interface and require zero or one step to revoke consent. According to Habib et al. [46], users who face persistent icons are more likely to recognise a correct method to revoke consent with respect to users who saw a link in the website’s footer. Therefore, even though all such implementations are compliant, only 7.4% websites with an Icon provide a more usable revocation design.
5.1.2 Two or more steps to revoke vs zero steps to accept
@@
33 (20.5%) websites out of 161 allow users to access the revocation option within the same interface, but with additional obstruction, requiring them 2 or more steps to revoke consent. Such websites (labeled with Options via Policy) hid the option to revoke consent inside the “Cookie Policy” or “Privacy Policy” page. According to the majority of EU regulators, this additional effort does not allow to exercise the right of revocation in a easy and effective way even if such option is located withing the interface. Arguably, the websites requiring additional effort neither comply with LR3 (Easy revocation through the same effort and number of steps) nor with the principles P1 (Fairness), since websites require unjustifiable and unexpected effort, nor P2 (Data Protection by Design), since the adopted measures are not efficient as to facilitate the revocation right but obstruct it instead, and P3 (Accountability), since such websites are not able to demonstrate compliance with the requirement LR3.
5.1.3 Revocation options Via Different Interface
@@
Overall, 32 (19.87%) websites out of 161 offered the option to revoke consent via an interface that is substantially different from the interface for users to accept consent. Figure 7 in the Appendix shows examples of websites for each of the options given.
@@
Two websites’ (tumblr.com and baidu.com) privacy settings lead the user to the login page and instruct the user to revoke consent after logging-in, which may require account creation (labeled After login). While this option occurs through the same website, it directs the user to a different interface. As before, this option infringes the requirement LR2 since the option redirects users to an unrelated interface. It also infringes principles P1 and P2 as it forces users to take unexpected additional steps and effort – to subscribe, and thus to give more personal data – to revoke consent. This option impedes these websites to show accountable revocation (P3).
5.1.4 No revocation option provided
Nine (5.6%) websites out of 161 did not provide any means to revoke consent, though they also did not display any consent banner, indicating that these websites probably did not integrate online tracking. We further analysed these websites and found that 4 (2.48%) of them stored AA cookies making them non-complaint. These websites include ntp.org and un.org. All these 4 websites that use AA cookies infringe the legal requirement LR1 (Right to revoke consent) and the following principles: P1 since websites process data unjustifiably for AA purposes without a legal basis, or the knowledge or consent of users; P2 as it does not create measures to enable the exercise of the revocation right. Consequently, these websites cannot be accountable for demonstrating compliance with this right (P3).
Figure 3: Revocation methods on top 161 websites
5.2 Impact of CMPs on compliance
@@
Regarding non-compliant implementation, 37% (39 out of 105) websites without CMP require two steps or more to revoke consent, while only 10.7% (6 out of 56) websites with CMP needed 2 or more steps for revocation. None of the websites with CMP denied a revocation option to users, though 8.5% (9 out of 105) websites without CMP offered no revocation option. In conclusion, websites with detected CMPs tend to be more compliant with revocation requirements.
6 Effect of Revocation on AA Cookies
@@
Figure 4: Change in the number of advertising and analytics (AA) cookies across different options
In this section, we address RQ2 and measure whether cookies requiring consent (Advertising and Analytics, or AA cookies) are deleted upon user’s consent revocation. As described in Section 4.1.2, we recorded the cookies of websites in four stages – upon initial landing, after accepting consent, after rejecting optional cookies, and after revocation. Out of the 120 websites where revocation was possible within the same interface, we find AA cookies after revocation on 69 (57.5%) websites.
Figure 4 compares the change in the number of AA cookies on websites after revocation w.r.t. other stages. Surprisingly, the number of AA cookies increases on most websites after revocation w.r.t. initial landing and w.r.t. after rejection (see red bar in Figures 4(a) and 4(b)). Additionally, on the majority of analysed websites, the number of AA cookies after acceptance remains the same after revocation (see pink bars in Figure 4(c)). Websites that add or keep AA cookies after consent revocation violate LR4 (Revoking requires stopping data processing and deletion of consent-based data), P1 and P2. Such AA cookies are processed unexpectedly and contrary to user’s decisions, without a legal basis and are thus illegal  [1, Art. 6(1)(a)].
7 Validity and Consistency of Consent
In this section, we analyse how consent information is stored and shared behind a website interface, whether it is consistently stored across different storage and APIs and whether it is legally valid, thus answering RQ3. We analysed 281 websites with detected CMPs (DS2 and DS3 from Table II) to check the validity and consistency of the consent string that these CMPs implement. Together with a legal expert co-author of the paper, we analyzed 11 purposes predefined within IAB Europe TCF v2.2 [35] (see Table IX in the Appendix) [^4] These purposes are largely consistent with those of the TCF v2.0 [37].. We determined that purposes 2-9 require consent and demand user action to be selected, thus such purposes cannot be enabled by default in TCString. Purposes 10-11 are exempted from this requirement and may be enabled by default. Purpose 1 is storage-based and enabled by default.
@@ 7 Validity and Consistency of Consent
For the consent strings extracted from OTAG variable, or from OptanonConsent cookie, we follow the OneTrust specification [70] and extract purpose numbers from its “groups” parameters. Since we do not have a precise specification of the meaning of these purposes, we cannot analyse which ones require consent. We therefore assume that a consent string contains a negative consent if it matches the value observed upon the initial landing stage, and a positive consent if the value contains more purposes than at initial landing.
@@
TABLE V: Websites where consent strings in browser storage and returned by APIs were positive in different phases.
7.1 Results
7.1.1 Positive consent upon initial landing, after rejection and after revocation
We first analyse whether the TCStrings returned by the __tcfapi can be considered as a “baseline” in our further comparison, that is whether such TCStrings are valid. For this purpose, we analysed whether the __tcfapi returns a positive consent at three stages: Initial landing, Rejection and Revocation, as shown in Table V.
After Rejection and Initial landing: Only two prior works [4, 40] analysed whether positive consent is stored or returned by APIs upon Initial landing and After Rejection. Compared to previous works, we found very few cases where positive consent is present after rejection in websites implementing the TCF (only 3 websites out of 150). In contrast, 10 websites, out of 164, store positive consent after rejection with OTAG.
@@ 7.1.1 Positive consent upon initial landing, after rejection and after revocation
7.1.2 Consent strings are often not updated after revocation
@@
Since a positive consent was present after revocation in numerous cases, we analysed whether consent strings were indeed properly updated when the user revoked consent on these websites (Table V). On websites implementing __tcfapi, we observe that 15 (9.3%) websites did not update the consent string, while 11 (10.9%) and 7 (21.9%) websites did not update the TCString stored in cookies, and in the localStorage respectively. Two websites that updated consent returned by __tcfapi are sourceforge.com and ft.com. Website sourceforge.com did not have a banner, stored a negative consent at initial landing, but updated it into a positive consent upon revocation. Website ft.com only updated the “consentScreen” parameter (which is unrelated to purposes) in the TCString after revocation, thus leaving a positive consent. Among websites implementing OTAG, 16 (10.5%) out of 152 did not update consent after user revocation. 6 websites have updated the consent string by adding or removing only some of the purposes, but the resulting consent string still contained a positive consent. Curiously, 3 out of 6 websites – cisco.com, opendns.com and webex.com – added one purpose to the consent string after revocation.
@@
7.1.3 Inconsistency among consent strings between browser storage and APIs
@@
Table VI shows the number of websites where the TCString found in the browser storage (cookies and localStorage) is not consistent with the TCString returned by the __tcfapi, and where the OptanonConsent cookie does not match the consent value returned by the OTAG. Such mismatches result in either an incorrect storage of consent or incorrect functioning of the APIs, and result in wrong consent being returned to third-parties. 5 websites implementing the TCF returned different TCStrings from __tcfapi and the TCString stored in the cookie after revocation. In 2 of such websites, (freep.com and megaphone.com) the TCString in the cookie was not updated, and still contained a positive consent, while the TCString returned by the __tcfapi was properly updated into a negative consent. In the website aol.com, while the __tcfapi returned a positive consent, the value in the cookie stores negative consent.
@@
In 3 websites (reuters.com, manchestereveningnews.co.uk and portfolio.com), the TCString returned by the __tcfapi showed a negative consent, while the TCString stored in the localStorage remained unchanged, i.e, it contained positive consent. On websites implementing OneTrust CMP, we observed 4 mismatches between the OTAG value and the OptanonConsent cookie value, where the cookies stored a negative consent and were shared on the network. All these mismatches can lead to the incorrect consent being shared to third-parties.
@@
These detected mismatches between consent strings across browser storage and APIs can be due to the fact that websites do not update, consistently, browser storage and APIs when users revoke consent. Consequently, websites do not correctly register consent, as they are obligated to, in order to assure that the registered consent is identical to the user’s choice in the user interface, thus, infringing the consent requirement LR5. Moreover, the principle P2 is not also complied with since websites did not implement technical measures to assure that revocation is done efficiently.
@@
TABLE VI: Inconsistent consent strings across browser storages and APIs after revocation. The numbers in brackets show the number websites that used both the storage and the API.
7.1.4 Inconsistency between consent returned via __tcfapi and consent shared on the network
@@
We investigate inconsistencies in the consent strings shared to and received from the third-parties, which, in turn, helps identify the responsible party for such inconsistencies. Previous works have only partially analysed such inconsistency [4, §VIII.A], where positive consent was found to be sent within a specific gdpr_consent URL parameter on websites where __tcfapi did not contain a positive consent string. We however observed that the consent can be shared via differently named URL parameter as well. Additionally, the consent string can be sent in the POST data as part of the request as well. We find different inconsistencies on 8 distinct websites out of the 136 websites where revocation is possible.
@@
These inconsistencies could arise due to three reasons. (1) Due to delay in updating the consent after user revokes consent; (2) due to some scripts not being updated about revocation in consent; (3) due to introducing/using a different TCString. We do not consider delay in sending the updated consent as a violation since it is implementation specific. However, there are 8 websites where we observe inconsistencies causing possible violations. Out of these 8 websites, 4 of them (forbes.com, time.com, n-tv.de and cadenaser.com) had a different TCString on the network. In the remaining websites (deadline.com, kotaku.com, manchestereveningnews.co.uk and walesonline.co.uk), we observed that the old positive consent string was shared on the network by some scripts even after revocation. Details regarding these inconsistencies can be found in Appendix .1.
@@
The detected mismatches between consent strings returned by __tcfapi and the TCString shared on the network (Table VIII) lead to an incorrect registration of consent revocation by websites and also by the implemented CMPs. Consequently, websites do not comply with LR5 since they do not correctly register user revocation, as they are obligated to, in order to assure that the registered consent is identical to the user’s choice in the user interface. As a result, a single TCString does not serve as a proof of revocation, which should consist of a negative consent. Moreover, the principle P2 is also not complied with since websites did not implement technical measures to assure that revocation is done efficiently.
8 Evaluating Correct Communication of Consent Revocation to Third-parties
@@
In this section, we address RQ4 and investigate if all the third-parties that were informed of the consent acceptance (either by accessing APIs or via HTTP requests) are also informed of consent revocation.
8.1 Results
8.1.1 Not all third-parties are informed of revocation by accessing consent via APIs
To examine which third-parties access the APIs implemented by CMPs, we override the implementation of __tcfapi and OTAG (see §4.2.4). Overall, we found 23 (9.6%) websites out of 238 where at least one third-party accessed the API to fetch positive consent after acceptance, but did not access the API to fetch revoked consent.
@@ 8.1.1 Not all third-parties are informed of revocation by accessing consent via APIs
Among these, on 163 websites, where __tcfapi is implemented (see Table V), on 14 (8.6%) websites at least one third-party requests the __tcfapi with the command “getTCData” after acceptance but not after revocation [^5] In the latest version of TCF, there is an option for third-parties to add an event-listener for consent update. However, since the implementation of this API is not standard, we could not intercept calls to it.. Out of 176 websites that support OTAG, on 13 (7.4%) websites, at least one third-party accesses the OTAG variable after accepting, but not after revocation. On 4 of these 13 websites, scripts from the third-party domain ads-static.conde.digital read the value of the variable, indicating that this variable can be used to know the consent choice given for AA purposes, since this domain is present in EasyList [72].
@@
While CMPs offer APIs and other methods for third-parties to access consent information, third-parties become responsible for requesting access to consent and for being informed when consent is revoked. Therefore the correct implementation of proper event listeners by the website or by CMPs becomes necessary to inform third-parties about consent revocation, specially since consent can be updated multiple times in a single user session.
@@
TABLE VII: Percentage of third-parties informed of acceptance but not informed of the revocation in 101 websites.
8.1.2 Not all third-parties are informed of the consent revocation via HTTP requests
On each website where consent revocation is possible, we detected third-parties that are informed of consent information via the TCStrings in the URL or postData (see §4.2.4). We then detect third-parties informed of consent after acceptance but not informed of revocation.
Of the 136 websites that implement __tcfapi and support revocation, 101 (74.2%) websites contain at least one third-party that was informed of consent after acceptance, but not after revocation via HTTP requests. Moreover, on 68 of these websites, third-parties that were not informed about revocation set cookies through the “Set-Cookie” HTTP header, and thus, did not stop processing user’s data after revocation. Table VII shows the percentage of third-parties not informed of consent after revocation. Surprisingly, 45 (44.5%) websites did not inform any of the detected third-parties, while 35 (34.6%) of them did not inform more than 75% of included third-parties. Figure 5 shows most prevalent third-parties that were not informed of the revoked consent via HTTP requests but were informed of consent acceptance. These websites include the big ad-tech industry actors doubleclick.net, criteo.com and adnxs.com. Interestingly, 73.77% of all third-party domains that do not receive a communication about consent revocation are present in EasyList [72] and therefore participate in advertising or tracking.
Figure 5: Top 15 third-parties not informed of the revoked consent via HTTP requests and the number of websites including these scripts but not informing them of revocation.
8.1.3 Failure to communicate consent revocation results in unlawful data processing

### v3

@@ 5.1.1. Compliant revocation interface (zero or one steps)
Out of the 158 websites, only 8 (5.6%) offer a persistent icon or button floating on the page, thus requiring zero steps to reach the revocation option. $6868$ (41.6%) websites offered a link option in the footer of the page, requiring one step to reach the revocation option. 6 (6.8%) websites showed a consent banner or icon when accessing the privacy policy page from the footer, therefore also requiring one step to revoke consent. All these implementations found on 82 (51%) websites (labeled as Icon, Footer Options and Banner on Policy in Table 4) comply with revocation GDPR principles and consent requirements (LR1-3) since they are presented within the same interface and require zero or one step to revoke consent. According to Habib et al. (Habib et al., 2022), users who face persistent icons are more likely to recognise a correct method to revoke consent with respect to users who saw a link in the website’s footer. Therefore, even though all such implementations are compliant, only 7.4% websites with an Icon provide a more usable revocation design.
5.1.2. Two or more steps to revoke vs zero steps to accept
@@
35 (22.1%) websites out of 158 allow users to access the revocation option within the same interface, but with additional obstruction, requiring them 2 or more steps to revoke consent. Such websites (labeled with Options via Policy) hid the option to revoke consent inside the “Cookie Policy” or “Privacy Policy” page. According to the majority of EU regulators, this additional effort does not allow to exercise the right of revocation in a easy and effective way even if such option is located withing the interface. Arguably, the websites requiring additional effort neither comply with LR3 (Easy revocation through the same effort and number of steps) nor with the principles P1 (Fairness), since websites require unjustifiable and unexpected effort, nor P2 (Data Protection by Design), since the adopted measures are not efficient as to facilitate the revocation right but obstruct it instead, and P3 (Accountability), since such websites are not able to demonstrate compliance with the requirement LR3.
5.1.3. Revocation options Via Different Interface
@@
Overall, 32 (19.87%) websites out of 158 offered the option to revoke consent via an interface that is substantially different from the interface for users to accept consent. Figure 7 in the Appendix shows examples of websites for each of the options given.
@@
On one website(tumblr.com) privacy settings lead the user to the login page and instruct them to revoke consent after logging-in, which may require account creation (labeled After login). While this option occurs through the same website, it directs the user to a different interface. As before, this option infringes the requirement LR2 since the option redirects users to an unrelated interface. It also infringes principles P1 and P2 as it forces users to take unexpected additional steps and effort – to subscribe, and thus to give more personal data – to revoke consent. This option impedes these websites to show accountable revocation (P3).
5.1.4. No revocation option provided
Nine (5.6%) websites out of 158 did not provide any means to revoke consent, though they also did not display any consent banner, indicating that these websites probably did not integrate online tracking. We further analysed these websites and found that 4 (2.48%) of them stored AA cookies making them non-complaint. These websites include ntp.org and un.org. All these 4 websites that use AA cookies infringe the legal requirement LR1 (Right to revoke consent) and the following principles: P1 since websites process data unjustifiably for AA purposes without a legal basis, or the knowledge or consent of users; P2 as it does not create measures to enable the exercise of the revocation right. Consequently, these websites cannot be accountable for demonstrating compliance with this right (P3).
Figure 3. Revocation methods on top 158 websites
5.2. Impact of CMPs on compliance
@@
Regarding non-compliant implementation, 37% (39 out of 105) websites without CMP require two steps or more to revoke consent, while only 10.7% (6 out of 56) websites with CMP needed 2 or more steps for revocation. None of the websites with CMP denied a revocation option to users, though 9 out of 105 websites without CMP offered no revocation option. In conclusion, websites with detected CMPs tend to be more compliant with revocation requirements.
6. Effect of Revocation on AA Cookies
@@
Figure 4. Change in the number of advertising and analytics (AA) cookies across different options
In this section, we address RQ2 and measure whether cookies requiring consent (Advertising and Analytics, or AA cookies) are deleted upon user’s consent revocation. As described in § 4.1.2, we recorded the cookies of websites in four stages – upon initial landing, after accepting consent, after rejecting optional cookies, and after revocation. Out of the 120 websites where revocation was possible within the same interface, we find AA cookies after revocation on 69 (57.5%) websites.
Figure 4 compares the change in the number of AA cookies on websites after revocation w.r.t. other stages. Surprisingly, the number of AA cookies increases on most websites after revocation w.r.t. initial landing and w.r.t. after rejection (see red bar in Figures 4(a) and 4(b)). Additionally, on the majority of analysed websites, the number of AA cookies after acceptance remains the same after revocation (see pink bars in Figure 4(c)). Websites that add or keep AA cookies after consent revocation violate LR4 (Revoking requires stopping data processing and deletion of consent-based data), P1 and P2. Such AA cookies are processed unexpectedly and contrary to user’s decisions, without a legal basis and are thus illegal  (European Parliament and Council of the European Union, 2016, Art. 6(1)(a)).
7. Validity and Consistency of Consent
In this section, we analyse how consent information is stored and shared behind a website interface, whether it is consistently stored across different storage and APIs and whether it is legally valid, thus answering RQ3. We analysed 281 websites with detected CMPs (DS2 and DS3 from Table 2) to check the validity and consistency of the consent string that these CMPs implement. Together with a legal expert co-author of the paper, we analyzed 11 purposes predefined within IAB Europe TCF v2.2 (IAB, 2025b) (see Table 9 in the Appendix). These purposes are largely consistent with those of the TCF v2.0 (Matte et al., 2020b). We determined that purposes 2-9 require consent and demand user action to be selected, thus such purposes cannot be enabled by default in TCString. Purposes 10-11 are exempted from this requirement and may be enabled by default. Purpose 1 is storage-based and enabled by default.
@@ 7. Validity and Consistency of Consent
For the consent strings extracted from OTAG variable, or from OptanonConsent cookie, we follow the OneTrust specification (OneTrust, [n. d.]) and extract purpose numbers from its “groups” parameters. Since we do not have a specification for the meaning of these purposes, we cannot analyse which ones require consent. We therefore assume that a consent string contains a negative consent if it matches the value observed upon the initial landing stage, and a positive consent if the value contains more purposes than at initial landing.
@@
Table 5. Websites where consent strings in browser storage and returned by APIs were positive in different phases.
7.1. Results
7.1.1. Positive consent upon initial landing, after rejection and after revocation
We first analyse if the TCStrings returned by __tcfapi can be considered as a “baseline” in our comparison, i.e., whether such TCStrings are valid. For this purpose, we analysed whether the __tcfapi returns a positive consent at three stages: Initial landing, Rejection and Revocation, as shown in Table 5.
After Rejection and Initial landing: Only two prior works (Matte et al., 2020a; Zhang et al., 2024) analysed whether positive consent is stored or returned by APIs upon Initial landing and After Rejection. Compared to previous works, we found very few cases where positive consent is present after rejection in websites implementing the TCF (only 3 websites out of 150). In contrast, 10 websites, out of 164, store positive consent after rejection with OTAG.
@@ 7.1.1. Positive consent upon initial landing, after rejection and after revocation
7.1.2. Consent strings are often not updated after revocation
@@
Since positive consent persisted post-revocation in many cases, we examined whether consent strings were properly updated (Table 5). Among websites using __tcfapi, 15 (9.3%) did not update the consent string, while 11 (10.9%) and 7 (21.9%) failed to update the TCString in cookies and localStorage, respectively. Notably, sourceforge.com lacked a banner, initially stored negative consent, but changed it to positive upon revocation. ft.com updated only the “consentScreen” parameter, keeping positive consent intact. For websites using OTAG, 16 (10.5%) out of 152 failed to update consent post-revocation. Additionally, six websites modified only some purposes while still retaining positive consent. Interestingly, cisco.com, opendns.com, and webex.com added one purpose to the consent string after revocation.
@@
7.1.3. Inconsistency among consent strings between browser storage and APIs
@@
Table 6 shows the number of websites where the TCString found in the cookies and localStorage is not consistent with the TCString returned by the __tcfapi, and where the OptanonConsent cookie does not match the consent value returned by the OTAG. Such mismatches result in either an incorrect storage of consent or incorrect functioning of the APIs, and result in wrong consent being returned to third-parties. 5 websites implementing the TCF returned different TCStrings from __tcfapi and the TCString stored in the cookie after revocation. In 2 of such websites, (freep.com and megaphone.com) the TCString in the cookie was not updated, and still contained a positive consent, while the TCString returned by the __tcfapi was properly updated into a negative consent. In the website aol.com, while the __tcfapi returned a positive consent, the value in the cookie stores negative consent.
@@
In three websites (reuters.com, manchestereveningnews.co.uk and portfolio.com), the TCString returned by __tcfapi showed negative consent, while the TCString stored in the localStorage remained unchanged, i.e, it contained positive consent. On websites implementing OneTrust, we observed 4 mismatches between OTAG and OptanonConsent cookie value, where the cookies stored a negative consent and were shared on the network. This mismatch can lead to the incorrect consent being shared to third-parties.
@@
These mismatches between consent strings across browser storage and APIs can be due to the fact that websites do not consistently update the storage and APIs when users revoke consent. Consequently, websites do not correctly register consent, as they are obligated to, thus, infringing the consent requirement LR5. Moreover, P2 is not also complied with since websites do not implement technical measures to assure that revocation is done efficiently.
@@
Table 6. Inconsistent consent strings across browser storages and APIs after revocation. The numbers in brackets show the number websites that used both the storage and the API.
7.1.4. Inconsistency between consent returned via __tcfapi and consent shared on the network
@@
We investigate inconsistencies in the consent strings shared to and received from the third-parties, which, in turn, helps identify the responsible party for such inconsistencies. Previous works have only partially analysed such inconsistency (Matte et al., 2020a, §VIII.A), where positive consent was found to be sent within a specific gdpr_consent URL parameter on websites where __tcfapi did not contain a positive consent string. We however observed that the consent can be shared via differently named URL parameter as well. Additionally, the consent string can be sent in the POST data as part of the request as well. We find different inconsistencies on 8 distinct websites out of the 136 websites where revocation is possible.
@@
These inconsistencies could arise due to: (1) delay in updating the consent after user revokes consent; (2) some scripts not being updated about revocation in consent; (3) introducing/using a different TCString. We do not consider delay in sending the updated consent as a violation since it is implementation specific. However, there are 8 websites where we observe inconsistencies causing possible violations. Out of these 8 websites, 4 of them (forbes.com, time.com, n-tv.de and cadenaser.com) had a different TCString on the network while deadline.com, kotaku.com, manchestereveningnews.co.uk and walesonline.co.uk, we observed that the old positive consent string was shared on the network by some scripts even after revocation. Details regarding these inconsistencies can be found in Appendix A.
@@
The detected mismatches between consent strings returned by __tcfapi and the TCString shared on the network (Table 8) lead to an incorrect registration of consent revocation by websites and also by the implemented CMPs. Consequently, websites do not comply with LR5 since they do not correctly register user revocation, as they are obligated to, in order to assure that the registered consent is identical to the user’s choice in the user interface. As a result, a single TCString does not serve as a proof of revocation, which should consist of a negative consent. Moreover, the principle P2 is also not complied with since websites did not implement technical measures to assure that revocation is done efficiently.
8. Communicating Consent Revocation to Third-parties
@@
Next, we address RQ4 and investigate if all the third-parties that were informed of the consent acceptance (either by accessing APIs or via HTTP requests) are also informed of consent revocation.
8.1. Results
8.1.1. Not all third-parties are informed of revocation by accessing consent via APIs
@@
To examine which third-parties access the APIs implemented by CMPs, we override the implementation of __tcfapi and OTAG (see §4.2.4). Overall, we found 23 (9.6%) websites out of 238 where at least one third-party accessed the API to fetch positive consent after acceptance, but did not access the API to fetch revoked consent. Among these, on 163 websites where __tcfapi is implemented (see Table 5), on 14 (8.6%) websites at least one third-party requests the __tcfapi with the command “getTCData” after acceptance but not after revocation [^3] In the latest version of TCF, there is an option for third-parties to add an event-listener for consent update. However, since the implementation of this API is not standard, we could not intercept calls to it.. Out of 176 websites that support OTAG, on 13 (7.4%) websites at least one third-party accesses the OTAG variable after accepting, but not after revocation. On 4 of these 13 websites, scripts from the third-party domain ads-static.conde.digital read the value of the variable, indicating that this variable can be used to know the consent choice given for AA purposes, since this domain is present in EasyList (Fanboy and Khrin, 2023).
@@
CMPs expose APIs to third-parties to access consent information making them responsible for requesting access to consent and for being informed on updates. Therefore, a proper implementation of event listeners by the website or by CMPs is necessary to inform third-parties about consent revocation, since consent can be updated multiple times in a single user session. We discuss about the standardization of the event-listeners in §9.
@@
Table 7. Percentage of third-parties informed of acceptance but not informed of the revocation in 101 websites.
8.1.2. Not all third-parties are informed of the consent revocation via HTTP requests
On each website where consent revocation is possible, we detected third-parties that are informed of consent information via the TCStrings and One Trust consent string in the URL or postData (see §4.2.4). We then detect third-parties informed of consent after acceptance but not informed of revocation.
Of the 136 websites that implement __tcfapi and support revocation, 101 (74.2%) websites contain at least one third-party that was informed of consent after acceptance, but not after revocation via HTTP requests. Moreover, on 68 of these websites, third-parties that were not informed about revocation set cookies through the “Set-Cookie” HTTP header, and thus, did not stop processing user’s data after revocation. Table 7 shows the percentage of third-parties not informed of consent after revocation. Surprisingly, 45 (44.5%) websites did not inform any of the detected third-parties, while 35 (34.6%) of them did not inform more than 75% of included third-parties. Figure 5 shows most prevalent third-parties that were not informed of the revoked consent via HTTP requests but were informed of consent acceptance. These websites include the big ad-tech industry actors doubleclick.net, criteo.com and adnxs.com. Interestingly, 73.77% of all third-party domains that do not receive a communication about consent revocation are present in EasyList (Fanboy and Khrin, 2023) and therefore participate in advertising or tracking.
Of the 152 websites that provide revocation and implement OneTrust, we observed that allowed categories from the OptanonConsent cookie (see §4.2.5) are rarely shared with third-parties, i.e., only on 8 websites OneTrust allowed categories were sent to third-parties after acceptance, and 6 sent the allowed categories after revocation. Interestingly, all third-party requests with allowed categories were made to Google Analytics (www.google-analytics.com) and Google Tag Manager (gtm.elementor.com). According to the official OneTrust documentation, the categories are communicated to the Google Tag Manager for it to manage consent and third parties (OneTrust, 2024)). We therefore conclude that OneTrust CMP may delegate the consent communication to the GTM framework and not send the allowed categories to third-parties by itself.
Figure 5. Top 15 third-parties not informed of the revoked consent via HTTP requests and the number of websites including these scripts but not informing them of revocation.
8.1.3. Failure to communicate consent revocation results in unlawful data processing